You deposit $65,000 in the bank today. What amount should you deposit in year 8, if you wish to accumulate $400,000 in year 20 if the interest rate is 6%?
Answer #1

Amount should be deposited in 8 years is $ 95,187.62

Step-1:Calculation of value of $ 65,000 in 20 years
Future value = P0 * (1+i)^n
= $     65,000.00 * (1+0.06)^20
= $     65,000.00 * 3.207135472
= $ 2,08,463.81
Where,
P0 = $     65,000.00
i = 6%
n = 20
Step-2:Calculation of remainig amount in 20 years
Tota required amount $ 4,00,000.00
Future value of above investment $ 2,08,463.81
Balance required future value $ 1,91,536.19
Step-3:Calculation of required deposit in 8 years
Required deposit in 8 years = FV * (1+i)^-n
= $ 1,91,536.19 * (1+0.06)^-12
= $ 1,91,536.19 * 0.496969364
= $ 95,187.62
Where,
FV = $ 1,91,536.19
i = 6%
n = 20-8 = 12